Grade 11Physical Chemistry

density of a gas at 276k and 380mm pressure is 1.5gm/litre. weight of 1.25 mole of the gas in grams is???

sreelekha
7 Years agoGrade 11

2 Answers

Arun
7 Years ago
M=dRT/PM=0.06793 kg1.23 M=.08492 kg=84.92 gm3PV=nRTPV=(m/M)RTPM=dRTP:Pressurem:mass of the gasM:Molecular Massd:DensityR:Universal Gas ConstantT:Temperature.P=380 mm= 0.5 atm=50662.5T=276 Kd=1.5g/l=1.5 kg/m
Arun
7 Years ago
PV=nRT PV=(m/M)RT PM=dRT P:Pressure m:mass of the gas M:Molecular Mass d:Density R:Universal Gas Constant T:Temperature. P=380 mm= 0.5 atm=50662.5 T=276 K d=1.5g/l=1.5 kg/m3 M=dRT/P M=0.06793 kg 1.23 M=.08492 kg=84.92 gm